How for an advance horizontal distance should t
Answer #1

height=200 m

speed=500m/s

let they stand x m from plane

time taken to travel 200 m down

200=0.5*9.8*t^2

t=6.39 sec

during this time horizontal distance travelled=x=500*6.39=3194.4 m

time taken=6.39 sec

horizontal component of velocity=500 m/s

vertical component=9.8*6.39=62.62 m/s downwards

plane will be just above the hikers when packet reaches the hiker as speed of plane is also 500 m/s while of packet is also 500 m/s horizontally